OpenStudy (anonymous):

1+x+x^2+...+x^n+... Find the interval of convergence

OpenStudy (anonymous):

is it -1<x<=1?

OpenStudy (anonymous):

\[\sum_{k=0}^{\infty}x^k=\frac{1}{1-x}\iff |x|<1\]

OpenStudy (anonymous):

strict inequality

OpenStudy (anonymous):

so lim(1^n) is divergent

OpenStudy (anonymous):

if x = 1 you get \[1+1+1+1+...\] not finite

OpenStudy (anonymous):

Using p series??????????

OpenStudy (anonymous):

and if x = -1 you get \[1-1+1-1+...\] which has no limit

OpenStudy (anonymous):

If the limit oscillate we have to further investigate it dude

OpenStudy (anonymous):

so why i have written this: \[\sum_{n=0}^{\infty}r ^{n}\] -1<r<=1 convergent all other r divergent

OpenStudy (anonymous):

the inequality on the right is wrong. it must be \[-1<r<1\] more succinctly written as \[|r|<1\]

OpenStudy (anonymous):

Matematika: the sequence r^n is a converging sequence when r=1, but the series isnt.

OpenStudy (anonymous):

you cannot sum up an infinite number of "1's" that is clear yes? \[1+1+1+...=\lim_{n\rightarrow \infty}n\]

OpenStudy (anonymous):

to be more precise, \[\sum_{k=0}^n1^n=n+1\] and by definition \[\sum_{n=0}^{\infty}1^n=\lim_{n\rightarrow \infty} \sum_{n=0}^{\infty}1^n=\lim_{n\rightarrow \infty}n+1=\infty\]

OpenStudy (anonymous):

typo in last line, should be \[\sum_{n=0}^{\infty}1^n=\lim_{n\rightarrow \infty} \sum_{n=0}^{n}1^n=\lim_{n\rightarrow \infty}n+1=\infty\]